# Comprehensive Review and Empirical Evaluation of Causal Discovery Algorithms for Numerical Data

### v1

@@ Abstract
Causal analysis has become an essential component in understanding the underlying causes of phenomena across various fields. Despite its significance, the existing literature on causal discovery algorithms is fragmented, with inconsistent methodologies and a lack of comprehensive evaluations. This study addresses these gaps by conducting an exhaustive review and empirical evaluation of causal discovery methods for numerical data, aiming to provide a clearer and more structured understanding of the field. Our research began with a comprehensive literature review spanning over a decade, revealing that existing surveys fall short in covering the vast array of causal discovery advancements. We meticulously analyzed over 200 scholarly articles to identify 24 distinct algorithms. This extensive analysis led to the development of a novel taxonomy tailored to the complexities of causal discovery, categorizing methods into six main types. Addressing the lack of comprehensive evaluations, our study conducts an extensive empirical assessment of more than 20 causal discovery algorithms on synthetic and real-world datasets. We categorize synthetic datasets based on size, linearity, and noise distribution, employing 5 evaluation metrics, and summarized the top-3 algorithm recommendations for different data scenarios. The recommendations have been validated on 2 real-world datasets. Our results highlight the significant impact of dataset characteristics on algorithm performance. Moreover, a metadata extraction strategy was developed to assist users in algorithm selection on unknown datasets. The accuracy of estimating metadata is higher than 80%. Based on these insights, we offer professional and practical recommendations to help users choose the most suitable causal discovery methods for their specific dataset needs.
@@ 1 Introduction
Causality Research, as a dynamically evolving interdisciplinary field, has garnered increasing scholarly attention. This section elaborates on the fundamental concepts, evolution, and classifications of causal analysis, addresses the limitations of previous studies, and introduces the innovative contributions and specific objectives of the current work.
@@
Causal analysis employs a systematic approach to uncovering the underlying causes of phenomena, primarily addressing the question of “Why.” Since Granger’s seminal work in 1969 (Granger, 1969), which introduced a mathematical concept of causality, the field has expanded from philosophy into economics (Imbens, 2004) and other domains such as medicine (Mani and Cooper, 2000), environmental science (Li et al., 2014), and dynamics (Hu et al., 2015). Recently, the rapid advancement of Artificial Intelligence (AI) has opened new avenues for causal analysis. The integration of machine learning has enhanced the precision and efficiency of data processing for causal inference, while causal learning has established a more reliable and trustworthy framework for machine learning (Scholkopf, 2019; Makhlouf et al., 2020). These two disciplines mutually reinforce each other, driving significant advancements in scientific research.
@@
Given that correlation does not imply causation, causal research necessitates a thorough investigation beyond simple association analysis. Pearl’s work has provided a widely accepted framework (Pearl et al., 2000), known as the “Ladder of Causation,” which delineates three stages: Association, Intervention, and Counterfactual. The initial stage, Association, involves observing relationships between variables, yet it is insufficient for identifying confounders or selection bias that may lead to spurious causation (Cheng et al., 2019). The second stage, Intervention, involves controlled experiments to quantify the causal impact of one variable on another. The final stage, Counterfactual analysis, requires a deep understanding of the causal mechanisms underlying the phenomena. Figure 1 shows the causation ladder and corresponding analysis engine (Bareinboim and Pearl, 2016).
Figure 1: Ladder of Causation
@@
Research on causal analysis is primarily categorized into two areas (Nogueira et al., 2022): Causal Inference and Causal Discovery. Causal Inference typically progresses from cause to effect, focusing on the quantitative problems of “Intervention Stage” within the causation ladder framework (Peters et al., 2017). The central concept of inference involves controlled trials (Yao et al., 2021), where both experimental and control groups are observed to determine the effects of interventions. In contrast, Causal Discovery seeks to identify causal relationships from observed outcomes, emphasizing the qualitative problems to learn causal structures (Gelman, 2011). Upon uncovering causal mechanisms, it becomes possible to infer outcomes based on hypothetical scenarios that have not occurred. Despite their inverse logical relationship, causal discovery and causal inference differ significantly in their research methodologies, algorithms, and applications.
@@
In practical scenarios, a significant challenge in causal analysis is managing diverse and complex data types. Standard data formats include time series (Eichler, 2012), cross-sectional, and panel data. Panel data, which has two dimensions (time and samples), combines elements of cross-sectional data, where all sample observations at a specific time point are included, and time series data, where observations of the same sample are recorded at different time points. Since time series data does not adhere to the assumptions of independent and identically distributed (i.i.d.) random variables, it necessitates specialized research on data processing techniques and causal analysis algorithms. Furthermore, due to the importance and widespread use of time series in real-world applications, many researchers have focused extensively on this data type (Assaad et al., 2021; Biswas and Mukherjee, 2024). If a time series can be viewed as a ”list”, then the i.i.d. variable is a ”set” without self-causes. Therefore, causal discovery for i.i.d. data, which differs from time-series causality, has also attracted a lot of research attention (Xie et al., 2019).
This study concentrates on two pivotal elements of causal analysis: (1) Causal Discovery and (2) Time Series and i.i.d. data analysis. An exhaustive investigation was conducted to delve into the methodologies of causal discovery, encompassing principles, algorithmic strategies, and recent advancements in the field.
Currently, the literature lacks a universally applicable algorithm for causal analysis (Edinburgh et al., 2021). Unlike causal inference, the precision of causal discovery heavily relies on the selection of an appropriate causality model. Hence, it is crucial to systematically organize and consolidate existing algorithms to effectively address various scenarios.
@@
Extensive efforts have been made to review and evaluate causal discovery algorithms. To address the challenge of limited benchmark datasets for causal discovery, many researchers have focused on developing data simulators in fields such as industrial systems (Menegozzo et al., 2022), neurology (Tu et al., 2019), and biology (Ma et al., 2023). Some works evaluated causal discovery methods through experiments, but they are not comprehensive as they only consider one type of methods. For example, Sogawa et al. (2010) evaluated the identification accuracy and robustness of linear non-Gaussian methods and its variants. Raghu et al. (2018) compared the performance of four conditional independence-based algorithms on mixed data with latent variables. Ko et al. (2018) summarized estimation of distribution algorithms (EDAs) and compared their performance on four datasets to infer the best ones.
@@
In addition, there are articles that survey various types of methods. Song et al. (2016) and Käding et al. (2021) compared causal discovery methods for bivariates on real-world bench datasets. However, their study did not include analysis of multivariate algorithms. Ombadi et al. (2020) evaluated four causal discovery algorithms on hydrometeorological data, aiming to guide researchers in determining which causal method is most appropriate based on the characteristics of hydrological system. Assaad et al. (2022) not only systematically organized time-series methodologies but also performed thorough evaluations of representative algorithms based on distinct causal structures.
@@
Unlike previous surveys, this project adopts a data-oriented approach, categorizing causal relationships into four types: i.i.d. causality, time-delay causality, instantaneous causality, and causal pairs. To address the research gap that often overlooks the user’s perspective, this study performs a comprehensive analysis starting from the intrinsic characteristics of the data (data assumptions). Motivated by this approach, comparative experiments were conducted, treating data assumptions as experimental factors and employing various algorithms as experimental subjects. These comparisons aim to establish the relationship between algorithms and specific data features within each causality category. Specifically, our pursuit is dedicated to identifying the optimal algorithm, considering various factors, including data size, linearity, stationarity, and noise attributes. By leveraging the extracted data features, users can choose the most appropriate algorithm for their specific needs.
The specific objectives have been enumerated to provide coherent guidance and tasks for this study:
@@
1. Systematic Investigation and Taxonomy of Causal Discovery Methods: Comprehensively collect and categorize methods and algorithms for causal discovery, summarizing the characteristics and applications of these algorithms.
2. Empirical Assessment of Algorithmic Performance through Comparative Experiments: Conduct horizontal comparisons of algorithms on datasets with different features and select appropriate evaluation metrics to assess their performance from multiple perspectives.
@@
3. Recommendation Mechanism for the Optimal Algorithm for Users: Provide meaningful insights and recommendations on the optimal algorithm for specific datasets, offering decision-making suggestions in various application fields.
The primary aim of this research is to develop recommendation algorithms tailored to specific data features through an exhaustive survey of existing methods and rigorous empirical analysis. Key contributions of this study include the creation of a method library featuring 24 algorithms for future experimental use and the execution of 16 comparative analyses to elucidate the relationship between algorithm performance and specific data features. The findings are extended to unknown datasets, highlighting their broader applicability and implications.
@@
The rest of this paper is organized as follows. Section 2 presents a comprehensive literature review, with Section 2.1 covering the fundamental preliminaries of the domain, Section 2.2 discussing assumptions for causality, Section 2.3 reviewing related works, Section 2.4 examining relevant surveys on causal analysis, and identifying research gaps in the current literature. Section 3 details the survey methodology, describing the method design in Section 3.1, the data collection process in Section 3.2, and the analysis of the survey data in Section 3.3. Section 4 focuses on causal discovery algorithms, beginning with Granger-based methods in Section 4.1, followed by condition independence-based methods in Section 4.2, state space dynamic-based methods in Section 4.3, structural equation modeling-based methods in Section 4.4, deep learning-based methods in Section 4.5, and hybrid methods in Section 4.6. Section 5 outlines the empirical study design, introducing the datasets used in Section 5.1, describing the evaluation metrics in Section 5.2, discussing the algorithms in Section 5.3, covering the environment settings in Section 5.4, and presenting the research questions guiding the study in Section 5.5. Section 6 presents the results analysis, addressing the research questions by comparing algorithm performance in Section 6.1, evaluating real-world applicability in Section 6.2, and exploring the generalization to unknown datasets in Section 6.3. Potential threats to validity are addressed in Section 7. Section 8 concludes this study, summarizing the main findings in Section 8.1 and offering recommendations for further work in Section 8.2.
2 Background
@@ 2.3 Related Works
In recent years, the topic of causal discovery has garnered significant interest as researchers endeavor to uncover causal relationships from observational and experimental data. To establish a solid foundation for understanding causal discovery, it is crucial to delve into the field’s core concepts and theories. An exhaustive search was conducted within the Web of Science repository to comprehensively encompass the pivotal papers in this domain, curating 301 relevant articles. In order to facilitate a more in-depth examination and exploration of these papers, a network diagram depicting co-cited literature (Figure 3) was generated.
The term “co-citation” means the phenomenon wherein two papers are simultaneously referenced by one or more subsequent papers, thereby establishing interlinked co-citation relationships between the two papers. In Figure 3, the threshold for the minimum number of co-cited articles was set at 15, affording a lucid means to observe literature of close thematic association, thereby furnishing a framework and cluster of causal discovery. We can see from the figure that foundational works by Judea Pearl, Peter Spirtes, and Clark Glymour have laid the groundwork for causal discovery.
Figure 3: Co-citation network
@@
The most prominent node, depicted in red, represents the seminal work of Peter Spirtes (2001). Pearl’s work on causal Bayesian networks and the introduction of causal graphical models significantly advanced the field’s theoretical foundation. These seminal contributions serve as the basis for further research and methodologies. Peter Spirtes introduced nonparametric structural causal models (SCM) (Pearl, 2009) as a formal and intelligible language for articulating causal knowledge and explaining causal notions used in scientific discourse. These include concepts like randomization, intervention, direct and indirect effects, confounding, counterfactuals, and attribution. The structural language’s algebraic component corresponds to the potential-outcome framework, while its graphical component incorporates Wright’s method of path diagrams. When combined, these components provide a robust approach for causal inference, addressing long-standing issues in empirical sciences, such as confounding control, policy evaluation, mediation analysis, and the algorithmization of counterfactuals.
Since the 1990s, conditional independence connections in data have been used to recover the underlying causal structure. One such constraint-based algorithm, developed by Spirtes and Glymour, is the PC algorithm. This algorithm starts with an undirected graph and recursively deletes edges based on conditional independence judgments. The PC algorithm assumes no confounders (unobserved direct common causes of two measured variables) and that the causal information retrieved is asymptotically true. However, the PC algorithm has limitations, especially with high-dimensional biological datasets, where its runtime grows exponentially with the number of variables, making it inefficient for large datasets. Additionally, the order of factors in the input data influences the output, reducing its reliability. Notably, Runge et al. (2019) proposed PCMCI, represented in a blue node in Figure 3, which extends the PC algorithm to time series data.
Another constraint-based algorithm is the FCI algorithm, also developed by Spirtes et al. The FCI algorithm starts with a fully connected undirected graph and removes edges connecting conditionally independent variables to create a causal graph. In the second phase, it orients edges by recognizing the “V” and “Y” structures and then orients the remaining edges using a set of criteria. FCI can produce asymptotically correct results even in the presence of confounders, making it more robust than PC in dealing with latent confounders. These methods can handle a wide range of data distributions and causal links and they have efficient means of conditional independence verification. It should be noted, however, that they may not always convey entire causal information because they yield equivalence classes that represent multiple causal structures satisfying the same conditional independencies (Spirtes and Zhang, 2016).
Score-based algorithms, which facilitate the identification of causal structures by optimizing a well-defined score function, do not entail any confounders. A well-known score-based method is Greedy Equivalence Search (GES) (Chickering, 2002), which directly searches the space of equivalence classes. GES iteratively adds or removes edges to improve the model’s score until no further modifications are possible. Another recent score-based technique is the Greedy Relaxation of the Sparsest Permutation (GRaSP) (Lam et al., 2022) algorithm developed by Lam et al. This algorithm aims to discover the most sparse permutation of variables that fulfills a specific set of causal constraints by greedily adding variables to the permutation while minimizing the impact on the sparsity measure.
Historically, deliberations on the causal relationship amid two factors have concentrated on the standard situation, where the factors included are Gaussian and have a Linear causal relationship. On the other hand, the discrete case has been considered. The Linear Gaussian circumstance, then again, is an exemption in which the causal asymmetry does not uncover itself in the observed data or their joint dispersion. In the event that the causal asymmetry exists, we should conclude that it must have been concealed somehow, either through latent factors or factual predisposition.
For causal discovery from continuous data, numerous techniques based on Functional Causal Models (FCMs) have recently been presented. The effect $Y$ is represented by an FCM as a function of the direct causes $X$ and certain unmeasurable components or noise.
Several variations of the FCM have been demonstrated to be capable of producing distinct causal directions and have found practical uses. The green node attributed to Shimizu (2006), introduces the linear, non-Gaussian, and acyclic model (LiNGAM). $f$ is linear in the LiNGAM and at most one of the noise terms and cause $X$ is Gaussian. The impact Y is further generated by a post nonlinear transformation on the nonlinear effect of the cause $X+$ noise term in the post nonlinear (PNL) causal model.
RCD (Repetitive Causal Discovery) is another causal functional model-based method recently developed by Maeda et. al. (2020) to uncover variable pairs influenced by latent confounders and determine the causal direction between variables unaffected by confounders. Experiments with simulated and real-world data are used to validate RCD’s effectiveness. RCD consists of three major phases. It begins by extracting a collection of ancestors for each variable, which indicate direct or indirect causes. The back-door criterion (Pearl et al., 2000) is used to infer causal directions between observable variables by taking into account the identified shared causes. The procedure is repeated until no additional changes occur. Second, RCD distinguishes direct causes from indirect causes via conditional independence inference. Finally, RCD finds correlated variable pairs with undetermined causal directions to identify pairs of variables impacted by the same latent confounders. As noted by Maeda et al., the experimental results show that RCD is effective at detecting latent confounders, determining causal directions, and detecting variable pairings that are affected by common confounders.
Spirtes and Zhang’s review (2016) comprehensively covers concepts and methodologies for causal inference, including manipulations, causal models, sample predictive modeling, causal predictive modeling, structural equation models, the causal Markov assumption, and the faithfulness assumption. Their review focuses on constraint-based causal structures, structural equation models, and mechanisms for distinguishing cause and effect.
Benozzo et al. (2016) in their recent studies propose a novel approach for causal discovery in time series. The developed technique is a classification- based causal discovery technique by defining a feature space fundamentally based on Granger Causality and by using the Multivariate Auto Regressive (MAR) model as the data generator. Cross Validation was used to assess the performance of the model and the results were quite promising. This makes it evident that this novel approach is a good substitute to the signal processing method for causality detection between two time series. However, as noted by Benozzo et. al., this approach is only limited to three time series at the moment.
Zheng et al. (2018) introduced DAG-NOTEARS in 2018, a method for causal discovery using continuous optimization. This method offers efficient global updates and avoids assumptions about the local graph structure, although it still faces challenges with nonconvex optimization problems. Another approach by Zheng et al., DAG-GNN, uses variational autoencoders with graph neural network architectures to address NP-hardness (Chickering, 1996; Chickering et al., 2004) in causal discovery.
Traditional approaches face NP-hardness due to the combinatorial nature of the task. NP-hardness is a common issue faced by traditional approaches for causal discovery. Zheng et al. proposes another method called DAG Structure Learning with Graph Neural Networks (DAG-GNN) to cater to this issue with the help of a continuous constraint that allows implementation of renowned continuous optimisation techniques by exploring the effectiveness of neural networks as functional approximators to develop a deep generative model. The method uses variational autoencoders parameterized with novel graph neural network architectures. The performance is tested over synthetic and real world datasets and the results achieved are promising.
Another recent advancement in this field was by Zhu et al. (2020) who introduced a novel approach to use Reinforcement Learning to search for the DAG with optimal score by incorporating a predefined score function and penalties for acyclicity. Zhu et al. use the actor-critic algorithm, implementing encoder-decoder models for building the actor. Post experimenting on synthetic and real datasets, the method gives promising results with 30-node graphs. However, dealing with larger graphs is still a challenge, as noted by Zhu et al. Moreover, there’s still scope for development of a more efficient core function and/or implementing other RL algorithms with appropriate early stoppage.
Recently, Wang et al. (2021) proposed a novel approach based on ordering based causal discovery with reinforcement learning. The ordering-based method for causal discovery used in this study is combined with reinforcement learning (RL), where an ordering is created and pruned using variable selection to engender a DAG. The two mechanisms are created using a Markov Decision Process (MDP) formulation along with an encoder-decoder architecture. The proposed approach has been tested on both artificial and real world datasets. Results show positive improvement as compared to other RL-based causal discovery approaches.
In this analysis, we compare and evaluate these novel approaches with conventional ones, considering recent advancements in the domain to provide a comprehensive evaluation of methods for causal discovery.
2.4 Relevant Survey on Causal Analysis
@@
In addition to the basic works and concepts mentioned in the previous section, we also need to investigate causal discovery from a more holistic perspective. Consequently, an exhaustive analysis must be conducted, combined with authoritative surveys over the past five years, to gain a comprehensive understanding of this field.

### v2

@@ Abstract
Causal analysis has become an essential component in understanding the underlying causes of phenomena across various fields. Despite its significance, existing literature on causal discovery algorithms is fragmented, with inconsistent methodologies, i.e., there is no universal classification standard for existing methods, and a lack of comprehensive evaluations, i.e., data characteristics are often ignored to be jointly analyzed when benchmarking algorithms. This study addresses these gaps by conducting an exhaustive review and empirical evaluation for causal discovery methods on numerical data, aiming to provide a clearer and more structured understanding of the field. Our research begins with a comprehensive literature review spanning over two decades, analyzing over 200 academic articles and identifying more than 40 representative algorithms. This extensive analysis leads to the development of a structured taxonomy tailored to the complexities of causal discovery, categorizing methods into six main types. To address the lack of comprehensive evaluations, our study conducts an extensive empirical assessment of 29 causal discovery algorithms on multiple synthetic and real-world datasets. We categorize synthetic datasets based on size, linearity, and noise distribution, employing five evaluation metrics, and summarize the top-3 algorithm recommendations, providing guidelines for users in various data scenarios. Our results highlight a significant impact of dataset characteristics on algorithm performance. Moreover, a metadata extraction strategy with an accuracy exceeding 80% is developed to assist users in algorithm selection on unknown datasets. Based on these insights, we offer professional and practical guidelines to help users choose the most suitable causal discovery methods for their specific dataset.
@@ 1 Introduction
Causality, as a dynamically evolving interdisciplinary field, has been gaining increasing attention from both academia and industry (Nogueira et al., 2021; Menegozzo et al., 2021; Masson-Delmotte et al., 2021; Ganguly et al., 2023). Causal analysis employs a systematic approach to uncovering the underlying causes of phenomena, primarily addressing the question of “Why” behind observed trends. Since Granger’s seminal work in 1969 (Granger, 1969), which introduced a mathematical concept of causality, the field has expanded from philosophy to economics (Imbens, 2004) and other domains such as medicine (Mani and Cooper, 2000), environmental science (Li et al., 2014), and dynamics (Hu et al., 2015). Recently, the rapid advancement of Artificial Intelligence (AI) has opened new avenues for causal analysis. The integration of machine learning has enhanced the precision and efficiency of data processing for causal inference, while causal learning has established a more reliable and trustworthy framework for machine learning (Scholkopf, 2019; Makhlouf et al., 2020). These two disciplines mutually reinforce each other, driving significant advancements in scientific research.
@@
Given that correlation does not imply causation, causal research necessitates a thorough investigation beyond simple association analysis. Pearl’s work (2000) provided a widely accepted framework, known as the “Ladder of Causation”, which delineates three stages: association, intervention, and counterfactual. The initial stage, association, involves observing relationships between variables, yet it is insufficient for identifying confounders or selection bias that may lead to spurious causation (Cheng et al., 2019). The second stage, intervention, involves controlled experiments to quantify the causal impact of one variable on another. The final stage, counterfactual analysis, requires a deep understanding of the causal mechanisms underlying the phenomena. Figure 1 shows the causation ladder and corresponding analysis engine (Bareinboim and Pearl, 2016).
Figure 1: Ladder of Causation (Runge et al., 2023a)
@@
This section elaborates on the fundamental concepts, evolution, and classifications of causal analysis, addresses the limitations of previous studies, and introduces the innovative contributions and specific objectives of the current work. Research on causal analysis is primarily categorized into two areas (Nogueira et al., 2022): causal inference and causal discovery. Causal inference typically progresses from cause to effect, focusing on the quantitative problems of “Intervention Stage” within the causation ladder framework (Peters et al., 2017). The central concept of inference involves controlled trials (Yao et al., 2021), where both experimental and control groups are observed to determine the effects of interventions. When the causal graph is known, observational data can be used to predict intervention effects of experimental tests. Common approaches for causal effect estimation include covariate adjustment (Pearl, 2009; Stekhoven et al., 2012; Maathuis and Colombo, 2015), optimal adjustment (Sekhon, 2008; Runge, 2021; Henckel et al., 2022), and the path method (Nandy et al., 2017). In contrast, causal discovery seeks to identify causal relationships from observed outcomes, emphasizing the qualitative problems to learn causal structures (Gelman, 2011). Upon uncovering causal mechanisms, it becomes possible to infer outcomes based on hypothetical scenarios that have not occurred. Despite their inverse logical relationship, causal discovery and causal inference differ significantly in their research methodologies, algorithms, and applications.
@@
In practical scenarios, a significant challenge in causal analysis is managing diverse and complex data types. Standard data formats include time series (Eichler, 2012), cross-sectional, and panel data. Panel data, which has two dimensions (time and samples), combines elements of cross-sectional data, where all sample observations at a specific time point are included, and time series data, where observations of the same sample are recorded at different time points. Since time series data does not adhere to the assumptions of independent and identically distributed (i.i.d.) random variables, it necessitates specialized research on data processing techniques and causal analysis algorithms. Furthermore, due to the importance and widespread use of time series in real-world applications, many researchers have focused extensively on this data type (Assaad et al., 2021; Biswas and Mukherjee, 2024). If a time series can be viewed as a “list", then the i.i.d. variable is a “set" without self-causes. Therefore, causal discovery for i.i.d. data, which differs from time-series causality, has also attracted a lot of research attention (Xie et al., 2019).
However, existing literature lacks a universally applicable algorithm for causal analysis (Edinburgh et al., 2021). Unlike causal inference, the precision of causal discovery heavily relies on the selection of an appropriate causality model. Hence, users often face challenges in selecting the appropriate causal discovery algorithm on unknown datasets, leading to unsatisfactory results or unnecessary expenditure of computational resources and time. The significance of this study including providing guidelines to help users quickly identify the most suitable algorithm for unknown datasets, and assisting researchers in organizing and benchmarking numerous existing algorithms. Therefore, we concentrates on two pivotal elements of causal analysis: (1) causal discovery and (2) time series and i.i.d. data analysis. An exhaustive investigation was conducted to delve into the methodologies of causal discovery, encompassing principles, algorithmic strategies, and recent advancements in the field.
@@
Extensive efforts have been made to review and reevaluate causal discovery algorithms in a unified, more extensive, and systematic way. To address the challenge of limited benchmark datasets for causal discovery, many researchers have focused on developing data simulators in fields such as industrial systems (Menegozzo et al., 2022), neurology (Tu et al., 2019), and biology (Ma et al., 2023). There have been some studies that conducted experimental studies to systematically evaluate one type of (not all existing) causal discovery methods. For example, Sogawa et al. (2010) evaluated the identification accuracy and robustness of linear non-Gaussian methods and its variants. Raghu et al. (2018) compared the performance of four conditional independence-based algorithms on mixed data with latent variables. Ko et al. (2018) summarized estimation of distribution algorithms (EDAs) and compared their performance on four datasets to infer the best ones.
@@
There have been articles that survey various types of methods. Song et al. (2016) and Käding et al. (2021) compared causal discovery methods for bivariates on real-world bench datasets. However, their research only focused on bivariate and did not investigate multivariate algorithms. Ombadi et al. (2020) evaluated four causal discovery algorithms on hydrometeorological data, aiming to guide researchers in determining which causal method is most appropriate based on the characteristics of hydrological system. Assaad et al. (2022) not only systematically organized time-series methodologies but also performed thorough evaluations of representative algorithms based on distinct causal structures.
@@
Unlike previous surveys, this paper adopts a data-oriented approach, categorizing causal relationships into four types: i.i.d. causality, time-delay causality, instantaneous causality, and causal pairs. To address the research gap that often overlooks the user’s perspective, this study performs a comprehensive analysis starting from the intrinsic characteristics of the data (data assumptions). Motivated by this approach, comparative experiments are conducted, treating data assumptions as experimental factors and employing various algorithms as experimental subjects. These comparisons aim to establish the relationship between algorithms and specific data features within each causality category. Specifically, our pursuit is dedicated to identifying the optimal algorithm, considering various factors, including data size, linearity, stationarity, and noise attributes. By leveraging the extracted data features, users can choose the most appropriate algorithm for their specific needs. The main contributions of this paper are summarized as follows:
  1. 1.
@@
    Survey and Taxonomy: We comprehensively collect and categorize methods and algorithms for causal discovery, summarizing the characteristics and applications of these algorithms.
  2. 2.
    Benchmarking: We conduct extensive benchmarking of selected state-of-the-art algorithms across diverse datasets using multiple evaluation metrics to access performance and applicability.
  3. 3.
@@
    Practical Guidelines: We provide practical insights and recommendations on the optimal algorithm for specific datasets, offering decision-making suggestions in various application fields.
In light of this, our experiment focus on the analysis of algorithm performance and result effectiveness, with the aim of addressing the following research questions:
This research question can be considered as a benchmark and baseline for answering other questions. By evaluating the impact of data features on algorithm performance, we establish a foundation that informs subsequent steps in our experimental analysis.
This research question is crucial for determining the effectiveness of insights gained from RQ 1, as it connects the findings from synthetic data to real-world scenarios. This question ensures that our conclusions are not limited to controlled experimental conditions but are also valid in practical applications.
RQ 3 (Generalization to unknown datasets): This RQ can be further divided into the following sub-research questions:
This research question is challenging with respect to experimental justifications of the other RQs as it involves extending the results of RQ 1 and RQ 2 to a broader range of applications. RQ 3.1 focuses on the feasibility of applying our findings to new datasets by analyzing their metadata, ensuring that our methods are robust and versatile. RQ 3.2 aims to translate these validated approaches into practical, user-friendly guidelines that assist practitioners in choosing the best algorithms for their unique datasets, thus bridging the gap between theoretical research and practical implementation.
Figure 2: The Overall Organization of the Study
@@
The rest of this paper is organized as follows. Section 2 presents the methodology of conducting the literature review, in which Section 2.1 covering the fundamental preliminaries of the domain, Section 2.2 discussing assumptions for causality, and Section 2.3 reviewing related algorithms and relevant surveys on causal analysis and identifying research gaps. Section 3 details the survey methodology, describing the data collection process in Section 3.1, and the analysis of the survey data in Section 3.2. Section 4 focuses on causal discovery algorithms, covering 6 categories of methodologies in Sections 4.1 to 4.6, respectively. Section 5 outlines the empirical study design, including investigated datasets in Section 5.1, evaluation metrics in Section 5.2, algorithms in Section 5.3, and environment settings in Section 5.4. Section 6 presents the results analysis, answering the research questions. Potential threats to validity are discussed in Section 7. Section 8 concludes this study and presents future work directions.
2 Background and Related Work
@@ 2.3 Related Works
Table 1: Timeline of Causal Discovery Algorithms
1969     Pairwise Granger Causality (Granger, 1969)one of the first statistical model of causal discovery
1982     Multivariate Granger Causality (MVGC) (Geweke, 1982; Chen et al., 2004; Barrett et al., 2010)
2001     PC (Spirtes et al., 2001; Colombo et al., 2014)turn observations into causal knowledge
2002     GES (Chickering, 2002a, b, 2020)
2006     ICALiNGAM (Shimizu et al., 2006)based on structural equation model
2008     ANM (Hoyer et al., 2008, 2012) Kernel Granger Causality (KGC) (Marinazzo et al., 2008, 2021) FCI (Zhang, 2008; Colombo et al., 2012)
2010     tsFCI (Entner and Hoyer, 2010) VARLiNGAM (Hyvärinen et al., 2010)
2011     IOTA (Hempel et al., 2011) DirectLiNGAM (Shimizu et al., 2011; Hyvärinen and Smith, 2013) ARMA-LiNGAM (Kawahara et al., 2011)
2012     CCM (Sugihara et al., 2012; Ye et al., 2015)dynamic systems PNL (Zhang and Hyvarinen, 2012) IGCI (Janzing et al., 2012)
2013     ES (Yuan and Malone, 2013) TiMINO (Peters et al., 2013)
2014     Copula Granger Causality (Copula GC) (Hu and Liang, 2014) CMS (Ma et al., 2014) PAI (McCracken and Weigel, 2014)
2015     oCSE (Sun et al., 2015)
2017     PSDR-TE (Mao and Shang, 2017)
2018     NOTEARS (Zheng et al., 2018, 2020; Fang et al., 2023) CGNN (Goudet et al., 2018) SCDA (Raghu et al., 2018) RECI (Blöbaum et al., 2018)
2019     PCMCI (Runge et al., 2019; Runge, 2020) TCDF (Nauta et al., 2019a) GraNDAG (Lachapelle et al., 2019) CDS (Fonollosa, 2019)
2020     CD-NOD (Huang et al., 2020) DYNOTEARS (Pamfil et al., 2020) RCD (Maeda and Shimizu, 2020; Maeda, 2022) GOLEM (Ng et al., 2020) NonSENS (Monti et al., 2020)
2021     CAM-UV (Maeda and Shimizu, 2021) DAG-GNN (Yu et al., 2019) CORL (Wang et al., 2021) NGC (Tank et al., 2021b; Wang et al., 2023)
2022     GRaSP (Lam et al., 2022) ACD (Löwe et al., 2022)
2024     NBCB, CBNB (Bystrova et al., 2024)
@@
In recent years, the topic of causal discovery has garnered significant interest as researchers endeavor to uncover causal relationships from observational and experimental data. To establish a solid foundation for understanding causal discovery, it is crucial to collect related works in this field. Pearl’s work (Pearl, 1985; Pearl et al., 2000) on causal Bayesian networks and the introduction of causal graphical models significantly advanced the field’s theoretical foundation. These seminal contributions serve as the basis for further research and methodologies. Peter Spirtes introduced nonparametric Structural Causal Models (SCM) (Pearl, 2009) as a formal and intelligible language for articulating causal knowledge and explaining causal notions used in scientific discourse. These include concepts like randomization, intervention, direct and indirect effects, confounding, counterfactuals, and attribution. The structural language’s algebraic component corresponds to the potential-outcome framework (Rubin, 1974), while its graphical component incorporates Wright’s method of path diagrams. The potential outcome framework, which focuses on estimating potential outcomes to calculate treatment effects, is particularly applicable to A/B tests. It performs effectively in causal inference, even when the complete causal graph is unknown (Aliprantis, 2015). When combined, these components provide a robust approach for causal inference, addressing long-standing issues in empirical sciences, such as confounding control, policy evaluation, mediation analysis, and the algorithmization of counterfactuals.
Table 1 illustrates the timeline of the development of causal discovery algorithms. As shown in Table 1, Granger (Granger, 1969) proposed a statistical model to determine causal relationships between bivariables, which became one of the oldest mathematical models in the history of causal discovery. Spirtes et al. introduced the assumptions and methods that laid the foundation for this field in their pioneering work (Spirtes et al., 2001), attempting to transform observations in real-world into causal knowledge. Spirtes and Glymour (2001) developed the PC algorithm as a fundamental constraint-based algorithm. This algorithm starts with an undirected graph and recursively deletes edges based on conditional independence judgments. Since 2006, Shimizu et al. (2006) have designed a Linear Non-Gaussian Acyclic Model (LiNGAM) algorithm based on Structural Equation Model (SEM), which was developed into many variants to address non-linear relationships, time series, mixed data, latent confounders, and other data cases, forming a class of algorithm groups. LiNGAM-related methods are widely applied across diverse fields, including neuroscience (Ji et al., 2024; Chiyohara et al., 2023), economics (Jin and Xu, 2024), epidemiology (Barrera and Miljkovic, 2022; García-Velázquez et al., 2020), psychology (Mojtabai, 2024; Rosenström et al., 2023), chemistry (Luo et al., 2024), and others. However, the above methods cannot handle nonseparable weakly connected dynamic systems. In response to this issue, Sugihara et al. (2012) proposed a Convergent Cross Mapping (CCM) algorithm based on state space method under the assumption of nonlinear deterministic systems, which is suitable for dynamic research fields such as ecology. In addition, for large sample datasets, causality algorithms combined with constantly evolving deep learning techniques greatly improve the accuracy of causal discovery and have become a popular research method.
@@
In addition to the algorithms and methods mentioned above, we also need to investigate causal discovery from a more holistic perspective. Consequently, an exhaustive analysis must be conducted, combined with authoritative surveys over the past five years, to gain a comprehensive understanding of this field.